# Visual Hand Gesture Recognition with Deep Learning: A Comprehensive Review of Methods, Datasets, Challenges and Future Research Directions

### v2

@@ 6.2 Pose-based
Several works exploit structured keypoint information by dividing the body into semantic regions such as the face, hands, and upper body. MSKA-SLR (MSKA_ELSEVIER) processes each region independently via attention-enhanced temporal convolutions across four branches. The resulting features are fused to decode gloss sequences using a CTC loss, complemented by frame-level self-distillation. Similarly, CoSign (CoSign) employs group-specific GCNs for each subset and introduces regularization to preserve inter-stream consistency. Their dual-stream variant integrates joint positions and motion features using multi-level fusion to capture both static and dynamic cues.
Temporal localization of gestures has also been explored in the literature. TY-Net (TY-Net) draws inspiration from YOLO to detect gesture instances in time by predicting their boundaries and centers, analogous to spatial bounding boxes.
Beyond joint coordinates, pose heatmaps serve as an alternative input modality. The multi-cue framework in (MULTI-CUE) combines spatial and temporal reasoning by localizing facial and hand regions via a pose estimator and subsequently encoding them with a temporal module, before decoding gloss sequences with a CTC head. The FA-STGCN model (FA-STGCN) further improves robustness to motion scale and frequency variability. It normalizes skeleton inputs and applies alternating spatial and temporal GCNs. Spatial modules learn latent joint relationships, while temporal processing is split into global and local paths—extracting coarse-grained and fine-grained motion patterns, respectively.
Finally, in the context of autonomous driving, Wang et al. (CTCX_WANG2022123) developed a two-stage pose-based system to classify road intersection gestures. The first stage extracts geometric features and encodes them via LSTM layers. The second stage uses a dual-stream setup, where one branch processes geometric descriptors and the other uses a ConvLSTM to handle co-occurrence matrices. Features are fused with directional orientation indicators for final classification.
@@ 6.3 Multimodal
A key example is TwoStream-SLR (NEURIPS2022_6cd3ac24_HEATMAPS), a dual-stream architecture for CSLR and SLT. It processes RGB and pose heatmaps in parallel via the first four S3D blocks in each branch, with bidirectional lateral connections for multi-level information exchange. Sign Pyramid Networks and auxiliary heads (video, keypoint, joint) provide additional supervision, while self-distillation minimizes KL-divergence across outputs. Fig. 8 illustrates the architecture. The same framework has been applied to ISLR (zuo2024onlinecontinuoussignlanguage) using a sliding window to extract sign boundaries and construct training dictionaries.
Figure 8: TwoStream-SLR (NEURIPS2022_6cd3ac24_HEATMAPS) is a dual-stream multimodal architecture built from eight S3D blocks. Bidirectional Lateral Connections enable multi-level information exchange between branches. Auxiliary classifiers, including the Sign Pyramid Network, Video Head, and Keypoint Head, provide additional supervision, while self-distillation transfers knowledge between the two streams.
Other CSLR models have adapted multimodal ISLR backbones. Uni-Sign (UNISIGNli2025) can be used directly for CSLR, while SignBERT+ (SignBERT+) can be extended by appending temporal pooling and a CTC decoder to generate gloss sequences.
Beyond sign language, general-purpose CHGR systems have explored multimodality across broader visual sensor inputs. For example, (pretrained_on_jester_3dDenseNet_LSTM) proposes a framework that segments RGB, depth, infrared, and optical flow inputs using a shared U-Net and processes them via separate 3D DenseNet-43 backbones. Extracted features are fused through Canonical Correlation Analysis to align shared representations across modalities.
@@ 6.4 Performance of CSLR Methods
Table 4 reports the computational cost and Word Error Rate (WER) of CSLR methods on three benchmark datasets. Leveraging large language models, as in (UNISIGNli2025) and (diffusion-CSLR-geng2026no), yields substantial performance gains without dramatically increasing parameters. SlowFastSign (SlowFastSign) achieves strong accuracy with a moderate number of parameters, while TwoStream-SLR (NEURIPS2022_6cd3ac24_HEATMAPS) has slightly higher WER, requires twice as many parameters, and involves a more complex training setup, though WER improves with additional training data (Fangyun_Wei_ICCV). C2ST (C2ST) is accurate but contains 187.8M parameters, which may limit real-world deployment. For resource-constrained scenarios, the pose-based CoSign-2s (CoSign) requires only 30.1 GFLOPs, though pose estimation cost is not included. DSTEN-CSLR (DSTEN-CSLR-yu2025dsten) is a lightweight RGB-based model but less accurate. Overall, a trade-off between accuracy and computational cost persists.
Table 4: Comparison of CSLR methods in terms of computational cost and accuracy (Word Error Rate, WER) on three benchmarks: CSL-Daily (Back-Translation-CVPR), Phoenix (PHOENIX-2014), and Phoenix T (NEURAL-SIGN-LANGUAGE-TRANSLATION-CVPR-CAMGOZ). ”Params” denotes the number of trainable parameters.
@@
(diffusion-CSLR-geng2026no) 2025 - 86.64M 23.7% 16.8% 17.2%
DSTEN-CSLR (DSTEN-CSLR-yu2025dsten) 2025 366.3 30.19M 28.1% 20.0% 20.3%
SLA-Adapter (CLIP_SLA_alyami2025clip_CSLR_2025) 2025 - 26.2M 25.8% 18.8% 19.5%
SLA-LoRA (CLIP_SLA_alyami2025clip_CSLR_2025) 2025 - 30.8M 25.8% 19.3% 19.4%
Uni-Sign (UNISIGNli2025) 2025 - - 26% - -
Swin-MSTP (SWIN_MSTP_Neurocomputing_2025_CSLR) 2025 - - 27.1% 18.7% 19.7%
CORE (CORE-CSLR-zhang2025core) 2025 - - 31.1% 21.7% 21.6%
SD-Net (SD-Net-CSLR-gao2025structure) 2025 - - 27.9% - 20.3%
STNet (STNet-wang2025continuous) 2025 - - 27.2% 19.3% 19.8%
SlowFastSign (SlowFastSign) 2024 - 52.5M 24.9% 18.3% 18.7%
MultiSignGraph (SignGraph_CVPR) 2024 - - 26.4% 19.1% 19.1%
CSGC (CSGC_Rao_Sun_Wang_Wang_Zhang_2024) 2024 - - 26.7% 19% 19.5%
CorrNet+ (corrnetplus_hu2024corrnetsignlanguagerecognition) 2024 - - 28.2% 18.2% 19.1%
TCNet (TCNet_Lu_Salah_Poppe_2024) 2024 932.1 19.3M 29.3% 18.9% 19.4%
C2SLR (C2SLR) 2024 1124.6 32.6M - 20.4% 20.4%
GPGN (GPGN) 2024 - - 30% 20.4% 20.5%
C2ST (C2ST) 2023 1635 187.8M 25.8% 17.7% 18.9%
CoSign-2s (CoSign) 2023 30.1 28.2M 27.2% 20.1% 20.1%
CTCA (CTCA) 2023 - - 29.4% 20.1% 20.3%
CorrNet (CorrNet) 2023 1035.4 20.5M 30.1% 19.4% 20.5%
SEN-CSLR (SEN-CSLR) 2023 1144.2 23.1M 30.7% 21% 20.7%
(Fangyun_Wei_ICCV) 2023 - 105.2M - 16.7% 18.5%
CVT-SLR (CVT-SLR) 2023 - - - 20.1% 20.3%
TwoStream-SLR (NEURIPS2022_6cd3ac24_HEATMAPS) 2022 - 105.2M 25.3% 18.8% 19.3%
STMC (MULTI-CUE) 2022 1742.1 40.71M - 20.7% 21%
TLP (TemporalLiftPooling) 2022 2950.5 48M - 20.8% 21.2%
SMKD (SMKD) 2021 1032.3 20.3M - 21% 22.4%
VAC-CSLR (VAC-CSLR) 2021 1120.4 22.3M - 22.3% -
@@ 7 Summary of SoTA Methods
Depth, pose, and other non-RGB modalities are used less frequently and typically in combination with RGB streams. This is because they either cannot represent fine-grained gestures (depth, optical flow) or perform poorly in dynamic recognition (pose estimators often fail on blurry images). RGB-based models—even older or simpler ones—already achieve strong results on SHGR datasets (HAGRIDv2_ARXIV), so adding pose can incur significant computational overhead without substantial accuracy gains. In CSLR, rapid hand movements degrade pose estimator performance, making pose usable only alongside RGB (NEURIPS2022_6cd3ac24_HEATMAPS) and increasing complexity. Thus, pose-based approaches are mostly limited to IDHGR, with some exceptions (CoSign), where incorporating multiple modalities can notably improve performance (SAM_SLR; jiang2021sign-sam-slr-v2), even with late fusion.
@@
Regarding architecture, SHGR’s simplicity allows even small 2DCNNs (EDenseNet; G-CNN-SHARMA2021115657) to achieve very high accuracy ($\geq$95%). However, the limited size and diversity of datasets like (NUS-I-2010; NUS-II-2013; ouhands; triesch2001system) can lead to overfitting, reducing real-world performance. In contrast, IDHGR and especially ISLR rely on more complex models. ST-GCN variants (JOINTS_MOTION_VECTORS_Improved_ST-GCN; TD-GCN) and GCN+Transformer pipelines (SignBERT_ICCV; SignBERT+; BEST; Weichao-Zhao) mainly process skeleton data, often combined with RGB 3DCNNs (e.g., I3D, Res3D) to compensate for pose estimation errors. While many studies focus on improving pose-based streams, RGB modalities are less explored. SAM-SLR (SAM_SLR) and SAM-SLRv2 (jiang2021sign-sam-slr-v2) propose advanced separate processing for each modality, representing current SoTA for IDHGR. RGBD models (RAAR3DNet; CTFB_Hampiholi; Yunan_Li) typically use dual-stream 3DCNNs, increasing computational cost. CSLR demands even more complex architectures to capture short-term (gloss-wise) and long-term (global) context. The 2DCNN+1DCNN+BiLSTM pipeline remains common, with relatively fewer parameters than SHGR models (e.g., Gesture-CNN (G-CNN-SHARMA2021115657)  67M vs. large CSLR models  48M (TemporalLiftPooling)). Vision transformers increasingly replace 2DCNN backbones (SWIN_MSTP_Neurocomputing_2025_CSLR; C2ST; CLIP_SLA_alyami2025clip_CSLR_2025) for their attention-based focus on critical regions. TwoStream-SLR (NEURIPS2022_6cd3ac24_HEATMAPS) employs a dual-stream 3DCNN with multimodal fusion and pose heatmaps (instead of skeleton joints) alongside RGB frames, achieving high accuracy.
@@
As mentioned earlier, SHGR is a relatively easy task, so researchers have not proposed any notable learning strategy. Pose-based methods for ISLR often (SignBERT_ICCV; SignBERT+; BEST; Weichao-Zhao) insert hand topology prior through SSL. NAS is also utilized to determine the optimal structure of 3DCNNs. Models for CSLR face serious optimization difficulties compared to other methods as a consequence of their complexity. Thus, extra supervision is provided via additional losses attached to various layers of the models (VAC-CSLR; CLIP_SLA_alyami2025clip_CSLR_2025). Leveraging additional training data combined with accurate architectures during training (Fangyun_Wei_ICCV) increases the performance, but makes a straightforward comparison with other methods infeasible.
@@
In summary, LHGR-Net (LHGR-Net) and stand-alone pose-based methods (Yangke_patcog; TD-GCN; CoSign) are suitable for resource-constrained use cases (assuming that the cost of the pose estimation is negligible, which is true especially when the keypoints are extracted by the deployed camera), balancing the computational complexity with the recognition accuracy. TwoStream-SLR (NEURIPS2022_6cd3ac24_HEATMAPS), C2ST (C2ST) and (diffusion-CSLR-geng2026no) achieve superior performance for CSLR but come with complex training settings and large computational cost. NLA-SLR (NLA_SLR) and SAM-SLRv2 (jiang2021sign-sam-slr-v2) surpass most of the related methods, but they require substantial computational resources. Uni-Sign (UNISIGNli2025) introduced a novelty, by combining various SLU tasks into a single one. Although it achieves significant performance and unifies a variety of tasks providing a shared framework, the computational complexity makes difficult the real-world deployment and the requirement of training data with detailed annotations prevents the adoption to other sign languages.
@@
FGDSNet (FGDSNet) SHGR HRI RGB Dual-stream hand segmentation model + dual-stream 2DCNN + mid-level fusion The dual-stream classifier addresses any failures of the task-specific segmentation model. The segmentation model and the 2DCNN backbones must be trained separately.
@@
LHGR-Net (LHGR-Net) SHGR HCI RGB Atrous convolutional layers followed by depth separable convolutions and attention. Suitable for resource-constrained deployment. Very small capacity.
@@
2DPSTPP-Net (2DPSTPP) SHGR HCI RGB 2DCNN with custom pooling layers and residual connections Spatial pyramid pooling and residual connections obtain robust representations. The model is relatively large for SHGR.
@@
Gesture-CNN (G-CNN-SHARMA2021115657) SHGR SLU RGB 2DCNN consisting of four convolutional layers The simplicity of the architecture enables easy implementation and training. Relatively large (67M parameters).
@@
SAM-SLR (SAM_SLR), SAM-SLR-v2 (jiang2021sign-sam-slr-v2) IDHGR SLU RGBD, Pose, HHA, OF, DF Six/seven streams followed by late fusion Incorporated a large number of modalities, achieving very high accuracy. Insufficient fusion and computationally expensive.
@@
SignBERT (SignBERT_ICCV), SignBERT+ (SignBERT+) (ISLR) IDHGR SLU RGB, Pose Spatial GCN + Transformer & 3DCNN + late fusion The model gains hand topology prior through SSL pretraining (masked hand tokens reconstruction, inspired by BERT) Needs extensive SSL pretraining, the processing and incorporation of the RGB modality may be suboptimal
@@
SignBERT+ (SignBERT+) (CSLR) CDHGR SLU RGB, Pose Spatial GCN + Transformer & 3DCNN + intermediate fusion + BiLSTM
@@
BEST (BEST) IDHGR SLU RGB, Pose dVAE + Spatial GCN + Transformer & 3DCNN + late fusion
@@
NLA-SLR (NLA_SLR) IDHGR SLU RGB, Pose Four-stream network based on S3D with two frame rates and multi-level fusion Novel architecture performing multimodality fusion efficiently. Language-aware label smoothing facilitates training. Robust to visually similar signs recognition. Is affected by the performance of word encoders
@@
Uni-Sign (UNISIGNli2025) IDHGR, CDHGR SLU RGB, Pose A pretrained LLM process features extracted by pose, visual and temporal encoders Provides a unified framework for multiple SLU tasks, increasing the recognition accuracy. Requires additional training data
@@
TD-GCN (TD-GCN) IDHGR - Pose ST-GCN variant Facilitates temporal modeling through non-fixed time-dependent topology Depends on the pose estimation performance.
@@
SBI-DHGR (SBI-DHGR) IDHGR - Pose Multi-stream 1DCNN + BiLSTM followed by late fusion Utilizes features of various levels of granularity. Very large compared to other pose-based methods
@@
DSTSA-GCN (DSTSA-CGN_cui2025dstsa) (4s) IDHGR - Pose Four stream GCN + TCN Group channel-wise and temporal-wise GCNs with dynamic graph topologies improve the performance. The static part of the hand topology is initialized randomly, lacking prior.
@@
SlowFast+NAS (3D_CDC_NAS_TIP_2021) IDHGR - RGBD, OF Multi-stream 3DCNN with 3D-CDC and multi-level fusion Enhanced temporal modeling through central difference convolution and optimal hyperparameter selection through NAS. Needs substantials computational resources (251.4M params, 232.1GFLOPs) compared to similar methods
@@
RAAR3DNet (RAAR3DNet) IDHGR - RGBD Dual stream I3D with adoptable cells, spatial and temporal attention NAS improves the architecture. The spatial attention mechanism is supervised by an off-the-shelf pose estimator. It relies on the correctness of the pose estimator
@@
CTFB (CTFB_Hampiholi) IDHGR - RGBD Dual-stream 3DCNN + Multi-level fusion with transformer-based modules Performs efficient multimodality fusion via multiple attention modules placed in different layers. The scaling to more than two modalities is not straightforward
@@
(diffusion-CSLR-geng2026no) CHGR SLU RGB Visual encoder followed by a Diffusion Model (DM). Using a DM for gloss generation makes it very accurate. Increased complexity.
@@
CLIP-SLA (CLIP_SLA_alyami2025clip_CSLR_2025) CHGR SLU RGB CLIP visual encoder (with MLP adopters or LoRA) + TCN + BiLSTM Early supervision by adding an extra classifier accelerates training, while the utilization of a foundation model enhances the feature extraction. The utilization of CLIP implies high computational cost.
@@
Swin-MSTP (SWIN_MSTP_Neurocomputing_2025_CSLR) CHGR SLU RGB SwinT+1DCNN/TLP Deploying a multiscale TCN, the model is able to recognize continuous signs of varying time durations. Computationally expensive.
@@
C2ST (C2ST) CHGR SLU RGB SwinT+1DCNN+BiLSTM CGD loss (that doesn’t make any independence assumption, compared to CTC) and self-distillation increase the performance. The vision transformer as well as the language model utilized during training increase the computational complexity.
@@
CoSign-2s (CoSign) CHGR SLU Pose ST-GCN variant The computational cost is negligible, thus CoSign-2s perfectly balances efficient execution and accurate recognition. The performance depends on the attributes of the pose estimator.
@@
(Fangyun_Wei_ICCV) CHGR SLU RGB, Pose TwoStream-SLR Utilization of extra training data via dictionary construction. Needs additional training data.
@@
TwoStream-SLR (NEURIPS2022_6cd3ac24_HEATMAPS) CHGR SLU RGB, Pose Dual-stream 3DCNN with bidirectional lateral connections Extra supervision by auxiliary losses and self-distillation applied to additional prediction heads facilitate training. The multi-level fusion incorporates efficiently the two modalities. Increased complexity due to the usage of two streams and an off-the-shelf pose estimator. Very large.

### v3

@@ 6.2 Pose-based
Several works exploit structured keypoint information by dividing the body into semantic regions such as the face, hands, and upper body. MSKA-SLR (62) processes each region independently via attention-enhanced temporal convolutions across four branches. The resulting features are fused to decode gloss sequences using a CTC loss, complemented by frame-level self-distillation. Similarly, CoSign (92) employs group-specific GCNs for each subset and introduces regularization to preserve inter-stream consistency. Their dual-stream variant integrates joint positions and motion features using multi-level fusion to capture both static and dynamic cues.
Temporal localization of gestures has also been explored in the literature. TY-Net (42) draws inspiration from YOLO to detect gesture instances in time by predicting their boundaries and centers, analogous to spatial bounding boxes.
Beyond joint coordinates, pose heatmaps serve as an alternative input modality. The multi-cue framework in (231) combines spatial and temporal reasoning by localizing facial and hand regions via a pose estimator and subsequently encoding them with a temporal module, before decoding gloss sequences with a CTC head. The FA-STGCN model (50) further improves robustness to motion scale and frequency variability. It normalizes skeleton inputs and applies alternating spatial and temporal GCNs. Spatial modules learn latent joint relationships, while temporal processing is split into global and local paths—extracting coarse-grained and fine-grained motion patterns, respectively.
Finally, in the context of autonomous driving, Wang et al. (204) developed a two-stage pose-based system to classify road intersection gestures. The first stage extracts geometric features and encodes them via LSTM layers. The second stage uses a dual-stream setup, where one branch processes geometric descriptors and the other uses a ConvLSTM to handle co-occurrence matrices. Features are fused with directional orientation indicators for final classification.
@@ 6.3 Multimodal
A key example is TwoStream-SLR (29), a dual-stream architecture for CSLR and SLT. It processes RGB and pose heatmaps in parallel via the first four S3D blocks in each branch, with bidirectional lateral connections for multi-level information exchange. Sign Pyramid Networks and auxiliary heads (video, keypoint, joint) provide additional supervision, while self-distillation minimizes KL-divergence across outputs. Fig. 8 illustrates the architecture. The same framework has been applied to ISLR (236) using a sliding window to extract sign boundaries and construct training dictionaries.
Figure 8: TwoStream-SLR (29) is a dual-stream multimodal architecture built from eight S3D blocks. Bidirectional Lateral Connections enable multi-level information exchange between branches. Auxiliary classifiers, including the Sign Pyramid Network, Video Head, and Keypoint Head, provide additional supervision, while self-distillation transfers knowledge between the two streams.
Other CSLR models have adapted multimodal ISLR backbones. Uni-Sign (113) can be used directly for CSLR, while SignBERT+ (77) can be extended by appending temporal pooling and a CTC decoder to generate gloss sequences.
Beyond sign language, general-purpose CHGR systems have explored multimodality across broader visual sensor inputs. For example, (125) proposes a framework that segments RGB, depth, infrared, and optical flow inputs using a shared U-Net and processes them via separate 3D DenseNet-43 backbones. Extracted features are fused through Canonical Correlation Analysis to align shared representations across modalities.
@@ 6.4 Performance of CSLR Methods
Table 4 reports the computational cost and Word Error Rate (WER) of CSLR methods on three benchmark datasets. Leveraging large language models, as in (113) and (59), yields substantial performance gains without dramatically increasing parameters. SlowFastSign (4) achieves strong accuracy with a moderate number of parameters, while TwoStream-SLR (29) has slightly higher WER, requires twice as many parameters, and involves a more complex training setup, though WER improves with additional training data (207). C2ST (217) is accurate but contains 187.8M parameters, which may limit real-world deployment. For resource-constrained scenarios, the pose-based CoSign-2s (92) requires only 30.1 GFLOPs, though pose estimation cost is not included. DSTEN-CSLR (214) is a lightweight RGB-based model but less accurate. Overall, a trade-off between accuracy and computational cost persists.
Table 4: Comparison of CSLR methods in terms of computational cost and accuracy (Word Error Rate, WER) on three benchmarks: CSL-Daily (230), Phoenix (49), and Phoenix T (27). ”Params” denotes the number of trainable parameters.
@@
(59) 2025 - 86.64M 23.7% 16.8% 17.2%
DSTEN-CSLR (214) 2025 366.3 30.19M 28.1% 20.0% 20.3%
SLA-Adapter (14) 2025 - 26.2M 25.8% 18.8% 19.5%
SLA-LoRA (14) 2025 - 30.8M 25.8% 19.3% 19.4%
Uni-Sign (113) 2025 - - 26% - -
Swin-MSTP (15) 2025 - - 27.1% 18.7% 19.7%
CORE (219) 2025 - - 31.1% 21.7% 21.6%
SD-Net (53) 2025 - - 27.9% - 20.3%
STNet (205) 2025 - - 27.2% 19.3% 19.8%
SlowFastSign (4) 2024 - 52.5M 24.9% 18.3% 18.7%
MultiSignGraph (52) 2024 - - 26.4% 19.1% 19.1%
CSGC (161) 2024 - - 26.7% 19% 19.5%
CorrNet+ (82) 2024 - - 28.2% 18.2% 19.1%
TCNet (124) 2024 932.1 19.3M 29.3% 18.9% 19.4%
C2SLR (234) 2024 1124.6 32.6M - 20.4% 20.4%
GPGN (64) 2024 - - 30% 20.4% 20.5%
C2ST (217) 2023 1635 187.8M 25.8% 17.7% 18.9%
CoSign-2s (92) 2023 30.1 28.2M 27.2% 20.1% 20.1%
CTCA (63) 2023 - - 29.4% 20.1% 20.3%
CorrNet (84) 2023 1035.4 20.5M 30.1% 19.4% 20.5%
SEN-CSLR (85) 2023 1144.2 23.1M 30.7% 21% 20.7%
(207) 2023 - 105.2M - 16.7% 18.5%
CVT-SLR (226) 2023 - - - 20.1% 20.3%
TwoStream-SLR (29) 2022 - 105.2M 25.3% 18.8% 19.3%
STMC (231) 2022 1742.1 40.71M - 20.7% 21%
TLP (83) 2022 2950.5 48M - 20.8% 21.2%
SMKD (69) 2021 1032.3 20.3M - 21% 22.4%
VAC-CSLR (136) 2021 1120.4 22.3M - 22.3% -
@@ 7 Summary of SoTA Methods
Depth, pose, and other non-RGB modalities are used less frequently and typically in combination with RGB streams. This is because they either cannot represent fine-grained gestures (depth, optical flow) or perform poorly in dynamic recognition (pose estimators often fail on blurry images). RGB-based models—even older or simpler ones—already achieve strong results on SHGR datasets (145), so adding pose can incur significant computational overhead without substantial accuracy gains. In CSLR, rapid hand movements degrade pose estimator performance, making pose usable only alongside RGB (29) and increasing complexity. Thus, pose-based approaches are mostly limited to IDHGR, with some exceptions (92), where incorporating multiple modalities can notably improve performance (91; 90), even with late fusion.
@@
Regarding architecture, SHGR’s simplicity allows even small 2DCNNs (189; 176) to achieve very high accuracy ($\geq$95%). However, the limited size and diversity of datasets like (106; 153; 132; 193) can lead to overfitting, reducing real-world performance. In contrast, IDHGR and especially ISLR rely on more complex models. ST-GCN variants (186; 120) and GCN+Transformer pipelines (78; 77; 224; 225) mainly process skeleton data, often combined with RGB 3DCNNs (e.g., I3D, Res3D) to compensate for pose estimation errors. While many studies focus on improving pose-based streams, RGB modalities are less explored. SAM-SLR (91) and SAM-SLRv2 (90) propose advanced separate processing for each modality, representing current SoTA for IDHGR. RGBD models (227; 66; 112) typically use dual-stream 3DCNNs, increasing computational cost. CSLR demands even more complex architectures to capture short-term (gloss-wise) and long-term (global) context. The 2DCNN+1DCNN+BiLSTM pipeline remains common, with relatively fewer parameters than SHGR models (e.g., Gesture-CNN (176)  67M vs. large CSLR models  48M (83)). Vision transformers increasingly replace 2DCNN backbones (15; 217; 14) for their attention-based focus on critical regions. TwoStream-SLR (29) employs a dual-stream 3DCNN with multimodal fusion and pose heatmaps (instead of skeleton joints) alongside RGB frames, achieving high accuracy.
@@
As mentioned earlier, SHGR is a relatively easy task, so researchers have not proposed any notable learning strategy. Pose-based methods for ISLR often (78; 77; 224; 225) insert hand topology prior through SSL. NAS is also utilized to determine the optimal structure of 3DCNNs. Models for CSLR face serious optimization difficulties compared to other methods as a consequence of their complexity. Thus, extra supervision is provided via additional losses attached to various layers of the models (136; 14). Leveraging additional training data combined with accurate architectures during training (207) increases the performance, but makes a straightforward comparison with other methods infeasible.
@@
In summary, LHGR-Net (218) and stand-alone pose-based methods (111; 120; 92) are suitable for resource-constrained use cases (assuming that the cost of the pose estimation is negligible, which is true especially when the keypoints are extracted by the deployed camera), balancing the computational complexity with the recognition accuracy. TwoStream-SLR (29), C2ST (217) and (59) achieve superior performance for CSLR but come with complex training settings and large computational cost. NLA-SLR (235) and SAM-SLRv2 (90) surpass most of the related methods, but they require substantial computational resources. Uni-Sign (113) introduced a novelty, by combining various SLU tasks into a single one. Although it achieves significant performance and unifies a variety of tasks providing a shared framework, the computational complexity makes difficult the real-world deployment and the requirement of training data with detailed annotations prevents the adoption to other sign languages.
@@
FGDSNet (229) SHGR HRI RGB Dual-stream hand segmentation model + dual-stream 2DCNN + mid-level fusion The dual-stream classifier addresses any failures of the task-specific segmentation model. The segmentation model and the 2DCNN backbones must be trained separately.
@@
LHGR-Net (218) SHGR HCI RGB Atrous convolutional layers followed by depth separable convolutions and attention. Suitable for resource-constrained deployment. Very small capacity.
@@
2DPSTPP-Net (89) SHGR HCI RGB 2DCNN with custom pooling layers and residual connections Spatial pyramid pooling and residual connections obtain robust representations. The model is relatively large for SHGR.
@@
Gesture-CNN (176) SHGR SLU RGB 2DCNN consisting of four convolutional layers The simplicity of the architecture enables easy implementation and training. Relatively large (67M parameters).
@@
SAM-SLR (91), SAM-SLR-v2 (90) IDHGR SLU RGBD, Pose, HHA, OF, DF Six/seven streams followed by late fusion Incorporated a large number of modalities, achieving very high accuracy. Insufficient fusion and computationally expensive.
@@
SignBERT (78), SignBERT+ (77) (ISLR) IDHGR SLU RGB, Pose Spatial GCN + Transformer & 3DCNN + late fusion The model gains hand topology prior through SSL pretraining (masked hand tokens reconstruction, inspired by BERT) Needs extensive SSL pretraining, the processing and incorporation of the RGB modality may be suboptimal
@@
SignBERT+ (77) (CSLR) CDHGR SLU RGB, Pose Spatial GCN + Transformer & 3DCNN + intermediate fusion + BiLSTM
@@
BEST (224) IDHGR SLU RGB, Pose dVAE + Spatial GCN + Transformer & 3DCNN + late fusion
@@
NLA-SLR (235) IDHGR SLU RGB, Pose Four-stream network based on S3D with two frame rates and multi-level fusion Novel architecture performing multimodality fusion efficiently. Language-aware label smoothing facilitates training. Robust to visually similar signs recognition. Is affected by the performance of word encoders
@@
Uni-Sign (113) IDHGR, CDHGR SLU RGB, Pose A pretrained LLM process features extracted by pose, visual and temporal encoders Provides a unified framework for multiple SLU tasks, increasing the recognition accuracy. Requires additional training data
@@
TD-GCN (120) IDHGR - Pose ST-GCN variant Facilitates temporal modeling through non-fixed time-dependent topology Depends on the pose estimation performance.
@@
SBI-DHGR (143) IDHGR - Pose Multi-stream 1DCNN + BiLSTM followed by late fusion Utilizes features of various levels of granularity. Very large compared to other pose-based methods
@@
DSTSA-GCN (32) (4s) IDHGR - Pose Four stream GCN + TCN Group channel-wise and temporal-wise GCNs with dynamic graph topologies improve the performance. The static part of the hand topology is initialized randomly, lacking prior.
@@
SlowFast+NAS (215) IDHGR - RGBD, OF Multi-stream 3DCNN with 3D-CDC and multi-level fusion Enhanced temporal modeling through central difference convolution and optimal hyperparameter selection through NAS. Needs substantials computational resources (251.4M params, 232.1GFLOPs) compared to similar methods
@@
RAAR3DNet (227) IDHGR - RGBD Dual stream I3D with adoptable cells, spatial and temporal attention NAS improves the architecture. The spatial attention mechanism is supervised by an off-the-shelf pose estimator. It relies on the correctness of the pose estimator
@@
CTFB (66) IDHGR - RGBD Dual-stream 3DCNN + Multi-level fusion with transformer-based modules Performs efficient multimodality fusion via multiple attention modules placed in different layers. The scaling to more than two modalities is not straightforward
@@
(59) CHGR SLU RGB Visual encoder followed by a Diffusion Model (DM). Using a DM for gloss generation makes it very accurate. Increased complexity.
@@
CLIP-SLA (14) CHGR SLU RGB CLIP visual encoder (with MLP adopters or LoRA) + TCN + BiLSTM Early supervision by adding an extra classifier accelerates training, while the utilization of a foundation model enhances the feature extraction. The utilization of CLIP implies high computational cost.
@@
Swin-MSTP (15) CHGR SLU RGB SwinT+1DCNN/TLP Deploying a multiscale TCN, the model is able to recognize continuous signs of varying time durations. Computationally expensive.
@@
C2ST (217) CHGR SLU RGB SwinT+1DCNN+BiLSTM CGD loss (that doesn’t make any independence assumption, compared to CTC) and self-distillation increase the performance. The vision transformer as well as the language model utilized during training increase the computational complexity.
@@
CoSign-2s (92) CHGR SLU Pose ST-GCN variant The computational cost is negligible, thus CoSign-2s perfectly balances efficient execution and accurate recognition. The performance depends on the attributes of the pose estimator.
@@
(207) CHGR SLU RGB, Pose TwoStream-SLR Utilization of extra training data via dictionary construction. Needs additional training data.
@@
TwoStream-SLR (29) CHGR SLU RGB, Pose Dual-stream 3DCNN with bidirectional lateral connections Extra supervision by auxiliary losses and self-distillation applied to additional prediction heads facilitate training. The multi-level fusion incorporates efficiently the two modalities. Increased complexity due to the usage of two streams and an off-the-shelf pose estimator. Very large.